# Birch SGD: A Tree Graph Framework for Local and Asynchronous SGD Methods

### v2

@@ E.4 Ringmaster ASGD
Algorithm 7 Ringmaster ASGD (Maranjyan et al.,, 2025)
1:  Input: point $w^{0}\in\mathbb{R}^{d}$, stepsize $\gamma>0,$ delay threshold $B\in\mathbb{N}$
@@
Condition 1 is satisfied because $\{\eta^{k-\delta^{k}}_{i}\}$ are i.i.d., $x^{k}=w^{k}$ and $z^{k}=x^{k-\delta^{k}}$ do not depend on $\xi^{k}=\eta^{k-\delta^{k}}_{i}.$ Condition 2 is satisfied because $\textnormal{repr}(z^{k})=\textnormal{repr}(w^{k-\delta^{k}})\subseteq\textnormal{repr}(w^{k})=\textnormal{repr}(x^{k}).$ Condition 3 is satisfied with $R=B-1$ because
$\textnormal{dist}(x^{k},z^{k})=\textnormal{dist}(x^{k},x^{k-\delta^{k}})=\delta^{k}\leq B-1,$
@@ E.6 Async-Local SGD
Similar to the previous proofs, Condition 1 is satisfied for the main branch $\{x^{k}\}_{k\geq 0}$ because all random variables $\{\eta^{i}_{j}\}$ in (17) are i.i.d., and $x^{0}$ and $z^{0}_{i_{1}}$ do not depend on $\eta^{0}_{i_{1}}.$ Points $x^{M-1}$ and $z^{M-1}_{i_{1}}$ do not depend on $\eta^{M-1}_{i_{1}},$ and so on. Conditions 2 is satisfied because all stochastic gradients used to compute $z^{p}_{i_{k}}$ are also used to compute the corresponding point on the main branch for all $p\in\{0,\dots,M-1\}$ and $k\geq 0$ (see Figure 5). Condition 3 is satisfied with $R=B-1+M-1=B+M-2$ due to the inequality $\delta^{k}=\textnormal{dist}(w^{k},z^{0}_{i_{k}})<B$ in Algorithm 9 and the fact every worker calculates $M$ stochastic gradients, which ensures that the tree distance between $z^{0}_{i_{k}}$ and the corresponding point from the main brain branch is at most $B-1,$ the tree distance between $z^{1}_{i_{k}}$ and the corresponding point from the main brain branch is at most $B-2,$ $\dots$,… , the tree distance between $z^{M-1}_{i_{k}}$ and the corresponding point from the main brain branch is at most $B+M-2.$
@@ E.9 Nested Local-Async SGD
Similarly to the previous proofs, Conditions 1 and 2 are satisfied by the construction of the algorithm. Using geometric graph reasoning, Condition 3 is satisfied with $R\leq B$ due to the requirement that “the total number of local steps in the clusters starting from the last broadcast is $B$.” This ensures that the distance between the points of the main branch and the corresponding points of the auxiliary sequence defined by $v^{\cdot}_{\cdot}$ does not exceed $B$. ∎
@@ E.10 Meta Local SGD
Compared to previous algorithms, Meta Local SGD is an abstract or meta-method, as it includes one abstract step: “Wait if needed and take any set of workers $S$.” This step is not explicitly defined, allowing users to apply any strategy they prefer. It may be random, where the algorithm chooses a uniformly random subset; it may follow a condition as in Local SGD, where the algorithm waits until $\sum_{i=1}^{n}M_{i}=B$; or it may be based on the current communication speeds of the workers, where the algorithm selects the workers with the fastest communication speeds at the current optimization moment.
@@ E.11 Dual-Process SGD
Initially, each worker waits for the first stochastic gradients with index 00 and immediately sends them once available. Then, while these are being transmitted, the workers continue their local computations. After the server receives the gradients with index 00, the workers begin sending the next batch of stochastic gradients, starting from index $1$ up to the latest index they have computed at that moment. This process continues until the server has received a total of $B$ stochastic gradients, accumulated through the communicated sums. This logic is implemented in Algorithm 19.
@@ Appendix F Computational Time Complexities of Algorithms under $h_{i}$-Fixed Computation Model
To compare methods, we consider the $h_{i}$-fixed computation model (Mishchenko et al.,, 2022). In this model, it is assumed that
@@
Note that it is possible to consider the universal computation model (Tyurin,, 2025) and capture virtually all possible computation behaviors of the workers. While the $h_{i}$-fixed computation model may seem more restrictive, it turns out that all optimal methods (Maranjyan et al.,, 2025) in the universal computation model are also optimal in the $h_{i}$-fixed computation model. Thus, for simplicity, we stick to the $h_{i}$-fixed computation model.
@@ F.2 Ringmaster ASGD
We use Lemma 4.1 from (Maranjyan et al.,, 2025).
@@
((Maranjyan et al.,, 2025)) Let the workers’ computation times satisfy the $h_{i}$-fixed computation model ((20) and (21)). Let $B$ be the delay threshold of Alg. 7. The time required to complete any $B$ consecutive iterate updates of Alg. 7 is at most
@@ F.6 Async-Local SGD
points (which is less than $B$ by assumption). This is because worker $i$ requires at most $Mh_{i}$ seconds to compute $M$ stochastic gradients before sending them to the server. Note that during any $B-1$ consecutive updates on the main branch, the server may ignore $M$ gradients from each worker at most once, because $\delta^{k}$ can be $\geq B$ at most once during $B-1$ consecutive updates. This explains the subtraction of $M$ in the formula.
@@ Appendix G Total Time Complexities of Algorithms under $(h,\tau)$-Fixed Computation Model
Note that it is possible to assume that each worker has its own communication time bound $\tau_{i}$ and computation time bound $h_{i}$ and consider $(h_{i},\tau_{i})$-fixed computation model (Tyurin et al.,, 2024). However, for simplicity, we assume $\tau_{i}=\tau$ and $h_{i}=h$ for all $i\in[n].$ See Section H for a more general case $(h_{i},\tau_{i})$-fixed computation model.
@@ G.1 Rennala SGD
(22)$=\mathcal{O}\left(h\times\left(\frac{L\Delta}{\varepsilon}+\frac{\sigma^{2}L\Delta}{n\varepsilon^{2}}\right)\right)$ )
@@ G.5 Ringmaster ASGD
assuming that the iteration rate $\Theta\left(\frac{BL\Delta}{\varepsilon}+\frac{\sigma^{2}L\Delta}{\varepsilon^{2}}\right)$ from Theorem E.4 is tight. As far as we know, this is the current state-of-the-art iteration rate of an Asynchronous SGD-like method (Maranjyan et al.,, 2025; Mishchenko et al.,, 2022; Koloskova et al.,, 2022; Cohen et al.,, 2021).
Appendix H Total Time Complexities of Algorithms under $(h_{i},\tau_{i})$-Fixed Computation Model
We now assume that each worker has its own communication time bound $\tau_{i}$ and computation time bound $h_{i}$ and consider the $(h_{i},\tau_{i})$-fixed computation model (Tyurin et al.,, 2024). It takes $\tau_{i}$ seconds to send a vector from worker $i$ to a parameter server and $\tau_{i}$ seconds to send a vector from the server to worker $i$ in the centralized setting. Alternatively, it takes $\tau_{i}$ seconds to send a vector to all other workers in the decentralized setting.
@@ Appendix H Total Time Complexities of Algorithms under $(h_{i},\tau_{i})$-Fixed Computation Model
Notice that Rennala SGD, Local SGD, and Cycle SGD wait for the slowest worker by the designs. If $\max_{i\in[n]}\tau_{i}\to\infty,$ then their total complexity tends to $\infty$.∞ . Thus, they are suboptimal under the $(h_{i},\tau_{i})$-fixed computation model. Ringmaster ASGD is suboptimal even under the $(h,\tau)$-fixed computation model. Async-Local SGD and Async-Batch SGD are optimal under the $(h,\tau)$-fixed computation model, but we conjecture that they are suboptimal under the $(h_{i},\tau_{i})$-fixed computation model.
We now prove that Dual-Process SGD is optimal under the $(h_{i},\tau_{i})$-fixed computation model.
H.1 Dual-Process SGD
Theorem H.1 (Dual-Process SGD).
@@ H.1 Dual-Process SGD
This complexity is optimal for methods without compression (Tyurin et al.,, 2024; Tyurin and Richtárik,, 2024). Notice that it is robust to slow communications. Indeed, if $\tau_{n}\to\infty,$ then this complexity will ignore worker $n$ due to the $\min\limits_{m\in[n]}$ operation.
Lemma H.2.
@@
Using Lemma H.2, the total time to find an $\varepsilon$–stationary point is
@@
Appendix I Performance of Rennala SGD and Ringmaster ASGD on a Quadratic Function

### v3

@@ E.4 Ringmaster ASGD
Algorithm 7 Ringmaster ASGD (Maranjyan et al., 2025)
1:Input: point $w^{0}\in\mathbb{R}^{d}$, stepsize $\gamma>0,$ delay threshold $G\in\mathbb{N}$
@@
$K\geq\frac{4GL\Delta}{\varepsilon}+\frac{8\sigma^{2}L\Delta}{\varepsilon^{2}}.$
with step size $\gamma=\min\{\frac{1}{2GL},\frac{\varepsilon}{4\sigma^{2}L}\}.$
@@
Condition 1 is satisfied because $\{\eta^{k-\delta^{k}}_{i}\}$ are i.i.d., $x^{k}=w^{k}$ and $z^{k}=x^{k-\delta^{k}}$ do not depend on $\xi^{k}=\eta^{k-\delta^{k}}_{i}.$ Condition 2 is satisfied because $\textnormal{repr}(z^{k})=\textnormal{repr}(w^{k-\delta^{k}})\subseteq\textnormal{repr}(w^{k})=\textnormal{repr}(x^{k}).$ Condition 3 is satisfied with R=G1R=G-1 because
$\textnormal{dist}(x^{k},z^{k})=\textnormal{dist}(x^{k},x^{k-\delta^{k}})=\delta^{k}\leq G-1,$
@@ E.6 Async-Local SGD
Similar to the previous proofs, Condition 1 is satisfied for the main branch $\{x^{k}\}_{k\geq 0}$ because all random variables $\{\eta^{i}_{j}\}$ in (17) are i.i.d., and $x^{0}$ and $z^{0}_{i_{1}}$ do not depend on $\eta^{0}_{i_{1}}.$ Points $x^{M-1}$ and $z^{M-1}_{i_{1}}$ do not depend on $\eta^{M-1}_{i_{1}},$ and so on. Conditions 2 is satisfied because all stochastic gradients used to compute $z^{p}_{i_{k}}$ are also used to compute the corresponding point on the main branch for all $p\in\{0,\dots,M-1\}$ and $k\geq 0$ (see Figure 5). Condition 3 is satisfied with R=B1+M1=B+M2R=B-1+M-1=B+M-2 due to the inequality $\delta^{k}=\textnormal{dist}(w^{k},z^{0}_{i_{k}})<B$ in Algorithm 9 and the fact every worker calculates $M$ stochastic gradients, which ensures that the tree distance between $z^{0}_{i_{k}}$ and the corresponding point from the main brain branch is at most B1,B-1, the tree distance between $z^{1}_{i_{k}}$ and the corresponding point from the main brain branch is at most B2,B-2, $\dots,$ the tree distance between $z^{M-1}_{i_{k}}$ and the corresponding point from the main brain branch is at most B+M2.B+M-2.
@@ E.9 Nested Local-Async SGD
Similarly to the previous proofs, Conditions 1 and 2 are satisfied by the construction of the algorithm. Using geometric graph reasoning, Condition 3 is satisfied with $R\leq B$ due to the requirement that “the total number of local steps in the clusters starting from the last broadcast is BB.” This ensures that the distance between the points of the main branch and the corresponding points of the auxiliary sequence defined by $v^{\cdot}_{\cdot}$ does not exceed $B$. ∎
@@ E.10 Meta Local SGD
Compared to previous algorithms, Meta Local SGD is an abstract or meta-method, as it includes one abstract step: “Wait if needed and take any set of workers SS.” This step is not explicitly defined, allowing users to apply any strategy they prefer. It may be random, where the algorithm chooses a uniformly random subset; it may follow a condition as in Local SGD, where the algorithm waits until $\sum_{i=1}^{n}M_{i}=B$; or it may be based on the current communication speeds of the workers, where the algorithm selects the workers with the fastest communication speeds at the current optimization moment.
@@ E.11 Dual-Process SGD
Initially, each worker waits for the first stochastic gradients with index 0 and immediately sends them once available. Then, while these are being transmitted, the workers continue their local computations. After the server receives the gradients with index 0, the workers begin sending the next batch of stochastic gradients, starting from index $1$ up to the latest index they have computed at that moment. This process continues until the server has received a total of $B$ stochastic gradients, accumulated through the communicated sums. This logic is implemented in Algorithm 19.
@@ Appendix F Computational Time Complexities of Algorithms under $h_{i}$-Fixed Computation Model
To compare methods, we consider the $h_{i}$-fixed computation model (Mishchenko et al., 2022). In this model, it is assumed that
@@
Note that it is possible to consider the universal computation model (Tyurin, 2025) and capture virtually all possible computation behaviors of the workers. While the $h_{i}$-fixed computation model may seem more restrictive, it turns out that all optimal methods (Maranjyan et al., 2025) in the universal computation model are also optimal in the $h_{i}$-fixed computation model. Thus, for simplicity, we stick to the $h_{i}$-fixed computation model.
@@ F.2 Ringmaster ASGD
We use Lemma 4.1 from (Maranjyan et al., 2025).
@@
((Maranjyan et al., 2025)) Let the workers’ computation times satisfy the $h_{i}$-fixed computation model ((20) and (21)). Let $B$ be the delay threshold of Alg. 7. The time required to complete any $B$ consecutive iterate updates of Alg. 7 is at most
@@ F.6 Async-Local SGD
points (which is less than $B$ by assumption). This is because worker $i$ requires at most $Mh_{i}$ seconds to compute $M$ stochastic gradients before sending them to the server. Note that during any B1B-1 consecutive updates on the main branch, the server may ignore $M$ gradients from each worker at most once, because $\delta^{k}$ can be $\geq B$ at most once during B1B-1 consecutive updates. This explains the subtraction of $M$ in the formula.
@@ Appendix G Total Time Complexities of Algorithms under $(h,\tau)$-Fixed Computation Model
Note that it is possible to assume that each worker has its own communication time bound $\tau_{i}$ and computation time bound $h_{i}$ and consider $(h_{i},\tau_{i})$-fixed computation model (Tyurin et al., 2024). However, for simplicity, we assume $\tau_{i}=\tau$ and $h_{i}=h$ for all $i\in[n].$ See Section I for a more general case $(h_{i},\tau_{i})$-fixed computation model.
@@ G.1 Rennala SGD
[…]
@@ G.5 Ringmaster ASGD
assuming that the iteration rate $\Theta\left(\frac{BL\Delta}{\varepsilon}+\frac{\sigma^{2}L\Delta}{\varepsilon^{2}}\right)$ from Theorem E.4 is tight. As far as we know, this is the current state-of-the-art iteration rate of an Asynchronous SGD-like method (Maranjyan et al., 2025; Mishchenko et al., 2022; Koloskova et al., 2022; Cohen et al., 2021).
Appendix H Comparison Between Our Local SGD and the Canonical Local SGD
In this section, we show that our version of Local SGD (Algorithm 5) achieves a better time complexity than the classical Local SGD. Although we focus in this section only on Local SGD, we expect similar improvements to extend to other new methods from Table 1. The purpose of this section is to highlight the tightness of the Birch SGD framework, using Local SGD as a case study.
In Section G.2, we prove that our version of Local SGD yields the total time complexity
$\Theta\left(\tau\frac{L\Delta}{\varepsilon}+h\left(\frac{L\Delta}{\varepsilon}+\frac{\sigma^{2}L\Delta}{n\varepsilon^{2}}\right)\right).$ (36)
We now illustrate that this result is provably better than the best theoretical result for the canonical version of Local SGD (Algorithm 20) known to us.
Algorithm 20 Local SGD (FedAvg) (McMahan et al., 2017)
0: initial model $x^{0}$, step size $\gamma$, # of local steps $K$
1:for $k=0,1,2,\ldots$ do
2:  Broadcast $x^{k}$ to all workers
3:  for each worker $i\in\{1,\ldots,n\}$ in parallel do
4:   $z^{k,0}_{i}=x^{k}$
5:   for $j=0,\ldots,K-1$ do
6:    $z^{k,j+1}_{i}=z^{k,j}_{i}-\gamma\nabla f(z^{k,j}_{i};\eta^{k,j}_{i})$
7:   end for
8:  end for
9:  $x^{k+1}=\frac{1}{n}\sum_{i=1}^{n}z^{k,K}_{i}$
10:end for
To the best of our knowledge, the state-of-the-art analysis of Algorithm 20 in the nonconvex setting is provided by Koloskova et al. (2020); Luo et al. (2025). Under Assumptions 1.1, 1.2, and 1.3, with a proper $\gamma,$ they establish the state-of-the-art iteration complexity
$\Theta\left(\frac{L\Delta}{\varepsilon}+\frac{L\sigma^{2}\Delta}{nK\varepsilon^{2}}+\frac{L\sigma\Delta}{K^{1/2}\varepsilon^{3/2}}\right)$
for finding an $\varepsilon$–stationary point for all $K\geq 1$. Next, under $(h,\tau)$-fixed computation model, this iteration complexity yields the time complexity
$\bar{T}:=\tau\left(\frac{L\Delta}{\varepsilon}+\frac{L\sigma^{2}\Delta}{nK\varepsilon^{2}}+\frac{L\sigma\Delta}{K^{\frac{1}{2}}\varepsilon^{\frac{3}{2}}}\right)+hK\left(\frac{L\Delta}{\varepsilon}+\frac{L\sigma^{2}\Delta}{nK\varepsilon^{2}}+\frac{L\sigma\Delta}{K^{\frac{1}{2}}\varepsilon^{\frac{3}{2}}}\right)$
(up to constant factors) because in each iteration the workers communicate, which takes $\tau$ seconds, and each worker (in parallel) computes $K$ stochastic gradients, which takes $h\times K$ seconds. Ignoring non-negative terms,
$\bar{T}\geq\tau\left(\frac{L\Delta}{\varepsilon}+\frac{L\sigma\Delta}{K^{\frac{1}{2}}\varepsilon^{\frac{3}{2}}}\right)+h\left(\frac{KL\Delta}{\varepsilon}+\frac{L\sigma^{2}\Delta}{n\varepsilon^{2}}+\frac{K^{\frac{1}{2}}L\sigma\Delta}{\varepsilon^{\frac{3}{2}}}\right)$
and $\bar{T}$ is lower bounded by
$\Theta\left(\sqrt{\tau h\frac{L^{2}\sigma^{2}\Delta^{2}}{\varepsilon^{3}}}+\tau\frac{L\Delta}{\varepsilon}+h\left(\frac{L\Delta}{\varepsilon}+\frac{\sigma^{2}L\Delta}{n\varepsilon^{2}}\right)\right)$ (37)
[…]
while
[…]
Note that the latter expression has a $\nicefrac{{1}}{{\varepsilon^{3/2}}}$ dependency, whereas our result has a $\nicefrac{{1}}{{\varepsilon}}$ dependency. Thus, our result is provably tighter.
Note that we obtain the time complexity (36) for several other new methods, including Async-Local SGD, Async-Batch SGD, and Dual-Process SGD.
Appendix I Total Time Complexities of Algorithms under $(h_{i},\tau_{i})$-Fixed Computation Model
We now assume that each worker has its own communication time bound $\tau_{i}$ and computation time bound $h_{i}$ and consider the $(h_{i},\tau_{i})$-fixed computation model (Tyurin et al., 2024). It takes $\tau_{i}$ seconds to send a vector from worker $i$ to a parameter server and $\tau_{i}$ seconds to send a vector from the server to worker $i$ in the centralized setting. Alternatively, it takes $\tau_{i}$ seconds to send a vector to all other workers in the decentralized setting.
@@ Appendix I Total Time Complexities of Algorithms under $(h_{i},\tau_{i})$-Fixed Computation Model
Notice that Rennala SGD, Local SGD, and Cycle SGD wait for the slowest worker by the designs. If $\max_{i\in[n]}\tau_{i}\to\infty,$ then their total complexity tends to $\infty.$ Thus, they are suboptimal under the $(h_{i},\tau_{i})$-fixed computation model. Ringmaster ASGD is suboptimal even under the $(h,\tau)$-fixed computation model. Async-Local SGD and Async-Batch SGD are optimal under the $(h,\tau)$-fixed computation model, but we conjecture that they are suboptimal under the $(h_{i},\tau_{i})$-fixed computation model.
We now prove that Dual-Process SGD is optimal under the $(h_{i},\tau_{i})$-fixed computation model within the family of methods that communicate either with a server (centralized setting) or with each other (decentralized setting).
I.1 Dual-Process SGD
Theorem I.1 (Dual-Process SGD).
@@ I.1 Dual-Process SGD
This complexity is optimal for distributed methods without compression communicating with a server (centralized setting) or with each other (decentralized setting) (Tyurin et al., 2024; Tyurin and Richtárik, 2024). Notice that it is robust to slow communications. Indeed, if $\tau_{n}\to\infty,$ then this complexity will ignore worker $n$ due to the $\min\limits_{m\in[n]}$ operation.
Lemma I.2.
@@
Using Lemma I.2, the total time to find an $\varepsilon$–stationary point is
@@
Appendix J Performance of Rennala SGD and Ringmaster ASGD on a Quadratic Function